FEL Upgrade Project Weekly Brief

April 1-5, 2002




This week we were pleased to welcome a visit by Virginia's new Secretary of Technology, George Newstrom and the Deputy Secretary, Eugene Huang on April 4th. Our guests were given a tour of CEBAF and the FEL Facility and participated in a ceremony honoring Jefferson Lab's recent patent award winners.


The preliminary agenda for the upcoming (May 1-3) semiannual review of the FEL Upgrade project was distributed this week to the review team and project stakeholders. Review preparations are well underway by the FEL Team.

We were pleased to receive the funding from ONR for the first half ($1.3M) of the efforts which were recently approved for FY02 activities in the amended ONR/DOE Memorandum of Agreement.

Discussions were held this week with Aerospace Corp. to revise our joint statement of work so that the US AF/DOE "Work for Others" could be amended for FY02 activities.


WBS 4 (Injector):

Continued working to support the W&M implantation system. A new support/mount for the support tube is being made along with 2 tubes for test runs. Polishing on the support tube and ball cathode continued. Detail part drawings for the new resistor are being worked.

Gun HVPS Tank - The last of the large transmission line piping was removed this week. This was in preparation to remove the old tank at the end of this month.

PARMELA simulations are underway regarding fine-tuning of the telescope quads to optimize

e-beam parameters at the entrance of the first LINAC.

A stainless steel piece one inch in diameter is being polished for further ion implantation. The effect of the implanted layer will be studied later under the Field Emission Scanning apparatus.




WBS 6 (RF):

Quarter HVPS - HV wiring in HVPS #3 is nearly complete, except for wiring to the klystron cart. Changed out all meters for HVPS #3, except for 2 that have not yet been received. The Control wiring for HVPS #3 is about 80% complete. The large capacitors were secured to the chassis of HVPS #3 this week and the missing MOV's were located.

Quarter Klystrons - The Machine Shop expects to start the plumbing work on the final 2 klystron carts. W. Lacy is helping them locate missing parts.

WBS 8 (Instrumentation):

The final cleanup off the cable trays took place this week. Many thanks to the I&C students and Jim! Next week the new cable installation will begin. The beam viewers will be followed by the trim magnets. Both of these systems have the new documentation complete. The software screens are nearly complete for the viewers, this was due to assistance from the CEBAF operations crews; thanks.

The fast raster power supplies upgrade has begun. There are minor modifications to the drive and control printed circuit board that need to be done. The markups are nearly complete, they will then enter the EECAD que. The design for the upgraded high voltage power supply control is also proceeding. This task will combine three chassis into one and present a much "cleaner" interface to the Machine Protection System (MPS).

The design and documentation for the Beam Loss Monitor VME conversion is nearly complete, we will go out for a prototype PCB next week. Front and rear panels are already in fabrication.

As a part of the tray cleanup the damaged ion pump HV cables were removed, we are awaiting the response from the cable manufacturer on resolving the cable failures. To prevent any electrical hazards the remaining high voltage power supplies were relocated to the tunnel so that NO active high voltage cables remain in the tray.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

At Magnet Enterprises International (MEI) in Oakland CA. The third coil looks very good. The fourth is being prepared for potting.

• Return legs remain in final machining. Top and bottom slabs are in grinding and 8 (for 4 magnets) are promised as a batch for April 12. Then the machining subcontractor will do final machining on the batch to create four magnet cores. There is now a possibility that the magnet will be complete by the Review on May 1.

Injector Dipoles (DU/DV)

• Master Machine is rough machining the slabs for the yokes.

• WANG NMR continues winding the GU and GV Coils.

Arc Dipoles (GY, GX, GQ)

• The contract for GQ-GX magnet cores was established with Process Equipment Co. of the Dayton Ohio area. They did all the yokes for the original CEBAF accelerator.

• The contract for the coils and assembly of the GQ-GX was established with Wang NMR. They are doing a very competent job at winding GY, GG, GU and GV coils.

• AES is magnetically modeling the GQ-GX dipoles at lower energy with shorter pole tips to match David Douglas’ use matrix. They will then design a shim system as part of Revision A for the pole tips to reconfigure the magnets for high energy use.

• Wang NMR of CA continues making the GY coil winding mandrel and designing and making the potting mold for the GY coils. They are also starting impregnation of the GG coils.

• Bosma Machine has the GY core pieces in final planning of the pole surface and the return legs in finish machining before grinding.


3 inch quad (QX)

• Some questions arose about the production measurements and hysteresis cycles matching the cycles in actual use and about the stability of the power supply being used. At week’s end, these questions are near resolution. In any case, there are no quadrupoles that are outside the specification for multipole content. Their slight differences in strength will be taken care of by individualized curves in the control system.

Trim Quad (QT)

• Milhous Control of Virginia reports that potting a first article was delayed. Potting didn't take place because a sample potting attempt caught a bad batch of epoxy that would not polymerize. A replacement batch is in transit.

Sextupole (SF)

• DULY Research performed additional optimization aimed at simplifying construction. At the same time, a construction method concept using quadrupole core machining technology is being incorporated in the mechanical design. Our thanks to New England Technicoil for discussions that led to this concept.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• Design of the ARC Chambers at AES continues, incorporating the new 100 kW aperture in the second arc GQ chambers. Overall design is very mature, with some chamber detail drawings at their second stage of review.

• The assembly and detail drawings for the girder for the region at the end of the cryomodules was signed.

• The assembly and detail drawings of the girders for the six quadrupole telescope in front of the first arc was completed, checked and is in final review at last.

• The Region in front of the first cryomodules including the injection line is laid out and detailed, ready for first review. The region includes stands and the X Chamber for the Injection dipoles.

WBS 11 (Optics):

The leadscrew and nut hardware are being procured. Design details continue on the vacuum vessels. Modeling of the 50kW laser beam dump is essentially complete for one concept (the flow channel design). Some results, such as the likelihood and impact of a decentered beam, need to be discussed and resolved. Design of the chamber and optics for the THz light pickoff continue.

The pump laser for the ultrafast laser system arrived. It will be installed in the OCR, and tested once the LSOP is modified and approved. The drive laser compressor hardware is installed in the drive laser enclosure and is being aligned before testing begins. Work continues on the OCMMS piezo transducer controls.

Other activities:

The Drive laser beam transport was partially dismantled to make room for gun HVPS activities. We assisted the Physics Div. with their Compton polarimeter rebuild.